# Strong discrete Morse theory

### v1

@@ 1. Introduction
Figure 1. Simple collapse (left) vs internal simple collapse (right).
@@ 1.1. Motivation and related work
A complementary perspective is presented in [26, 25], where the authors encode the homotopy type of a simplicial complex endowed with a discrete Morse function into a combinatorially defined flow category. The classifying space of the flow category recovers the homotopy type of the original complex. However, these categories are often large and computationally challenging, limiting their algorithmic usage in homotopy reconstruction.
@@ 1.2. Main results and outline
The combinatorial structure of the strong internal core admits a computational interpretation of internal strong collapses, which in turn leads to an efficient algorithm for their computation (Algorithm 2). An implementation is publicly available at https://github.com/ximenafernandez/Strong-Morse-Theory. To illustrate the effectiveness of this method, we apply it to examples from the Library of Triangulations [6], a repository of challenging simplicial complexes frequently used in the study of homotopy types. The results of the reductions are presented in Tables 1 and 2.
@@
The remainder of this manuscript is organized as follows. Section 2 describes discrete Morse theory as generalized Whitehead’s collapses. In Section 3, we study a strong version of discrete Morse theory as a generalization of strong collapses, which allows us to explicitly recover the combinatorial structure of the reduced CW-complex. Section 4 describes an algorithm for the construction of the strong internal core in terms of posets and acyclic matchings. Finally, Section 5 discusses the algorithm for random strong Morse reduction and presents experiments on the Library of Triangulations [6].
@@ 2. Discrete Morse theory as generalized collapses
We denote by ${\mathrm{core}}_{f}(K)$ the CW-complex obtained from the critical simplices of $f$ in Theorem 2.3, and refer to it as the internal core of $K$ associated to $f$. It is important to note that Theorem 2.3 does not provide an explicit description of how the critical cells are attached along their boundaries in ${\mathrm{core}}_{f}(K)$. In [15, Prop. 1.3, Thm. 1.6], the author offers a recursive construction of the attaching maps for the critical cells, based on the finiteness of $K$; however, the construction is only made explicit for 2-dimensional cells.
@@
Proposition 2.5 also admits a formulation in terms of CW-complexes; see [15, Proposition 1.3]. We refer to the deformation from $K$ to $L$ as an internal collapse. The collapse takes place within a subcomplex $K_{0}\subseteq K$, while the resulting deformation affects the entire complex through the reattachment of cells corresponding to the simplices in $K\smallsetminus K_{0}$.
@@ 3. Strong discrete Morse theory
Given a simplicial complex $K$ and a vertex $v\in V(K)$, the open star of $v$, denoted ${\mathrm{st}}(v,K)$, is the collection of simplices in $K$ that contain $v$. The subcomplex of simplices disjoint from $v$ is denoted $K\smallsetminus v$. The link of $v$ in $K$, denoted ${\mathrm{lk}}(v,K)$, is the subcomplex of $K\smallsetminus v$ consisting of simplices $\sigma$ such that $\sigma\cup\{v\}$ is a simplex of $K$. A simplicial cone over $K$ with apex $a$ (a vertex not in $K$) is the complex $a*K$ consisting of all simplices of $K$, the vertex $\{a\}$, and all simplices of the form $\sigma\cup\{a\}$ with $\sigma\in K$.
@@
Let $K$ be a simplicial complex. A vertex $v$ of $K$ is dominated by a vertex $a$ if ${\mathrm{lk}}(v,K)$ is a simplicial cone $a*K_{0}$, where $K_{0}$ is a subcomplex of $K$. In that case, there is an elementary strong collapse from $K$ to $K\smallsetminus v$, denoted $K\searrow K\smallsetminus v$. A strong collapse $K\searrow\!\searrow L$ from a simplicial complex $K$ to a subcomplex $L$ is a sequence of elementary strong collapses starting at $K$ and ending at $L$. In that case, we say that $L$ is a strong core of $K$. If $L$ has no dominated vertices, we say that $L$ is a minimal strong core of $K$ [^2] This is simply called a core in [3]. We will use the term minimal strong core in this manuscript to avoid confusion with other core constructions.
If $v\in K$ is dominated by $a$, then there is an explicit strong deformation retraction $r_{a}\colon K\to K\smallsetminus v$ given by
$r_{a}(\sigma)=\begin{cases}\sigma&\text{if }v\notin\sigma,\\ \{a\}\cup\sigma\smallsetminus\{v\}&\text{if }v\in\sigma.\end{cases}$
@@
with $K_{i+1}=K_{i}\smallsetminus v_{i}$ and $v_{i}$ dominated by $a_{i}\in K_{i}$, then there is a strong deformation retraction $r\colon K\to L$ given by the composition
$r:=r_{a_{n}}\circ r_{a_{n-1}}\circ\dots\circ r_{a_{1}}.$
@@
Recall that a CW-complex is regular if the attaching maps of all its cells are homeomorphisms onto their images. The following lemma formalizes the notion of an internal strong collapse, illustrated in Figure 2.
Lemma 3.3.
Let $K$ be a simplicial complex, and let $w$ be a vertex of $K$. Suppose there exists a subcomplex $L\subseteq K\smallsetminus w$ such that $K\smallsetminus w{\searrow\searrow\hskip 3.0pt}L$. Then $K$ is homotopy equivalent to a regular CW-complex
$Z:=L\cup\bigcup_{i}e_{i},$
where $L$ is embedded as a subcomplex of $Z$, and for each $k$-simplex $\sigma_{i}\in{\mathrm{st}}(w,K)$, there exists a corresponding $k$-cell $e_{i}$ not in $L$. The attaching map of each cell $e_{i}$ is induced from the attaching map of $\sigma_{i}$ and the strong deformation retraction $r\colon K\smallsetminus w\to L$ induced by the strong collapse.
If $\dim(K)=N$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0ptZ$.
@@
where $i$ is the inclusion and $r\colon K\smallsetminus w\to L$ is a strong deformation retraction induced by a strong collapse. By the universal property of the pushout, there exist induced maps $\bar{r}\colon K\to Z$ and $\bar{i}\colon L\to Z$ such that $\bar{r}\circ i=\bar{i}\circ r$. Since $r$ is a homotopy equivalence and $i$ is a closed cofibration, the Gluing Theorem (see, e.g., [11, 7.5.7, Corollary 2]) implies that $\bar{r}$ is also a homotopy equivalence.
The space $Z$ can be identified with the adjunction space $(K\sqcup L)/_{\sim}$, where $i(\sigma)\sim r(\sigma)$ for every simplex $\sigma\in K\smallsetminus w$. Let $q\colon K\sqcup L\to Z$ denote the corresponding quotient map. We define a CW-structure on $Z$ as follows. For each simplex $\sigma\in L$, we have $r(\sigma)=\sigma$, so $q$ acts as the identity and embeds $L$ into $Z$ with the same CW-structure it inherits from $K$. Moreover, since $r$ is simplicial, each cell $q(\sigma)$ for $\sigma\in K\smallsetminus w$ inherits a simplicial (hence regular) CW-structure from $K$. Now let $\sigma\in\operatorname{st}(w,K)$, and assume first that the strong collapse $K\smallsetminus w\searrow L$ is elementary, i.e., $L=(K\smallsetminus w)\smallsetminus v$ for some vertex $v$ dominated by $a\in K\smallsetminus w$. If $v\notin\sigma$, then $q(\sigma)=\sigma$ in $Z$, and the simplicial CW-structure is preserved. Similarly, if $v\in\sigma$ but $a\notin\sigma$, then $q(\sigma)=\{a\}\cup(\sigma\smallsetminus\{v\})$, which corresponds to a relabelling and again preserves regularity. The remaining case is when both $v$ and $a$ belong to $\sigma$. Suppose $\sigma=\{w,a,v_{0},\dots,v_{k}\}$, with $v_{0}=v$. Let $\bar{\sigma}\subseteq|K|$ denote the geometric realization of $\sigma$, which consists of formal convex combinations
$s\cdot a+r\cdot w+\sum_{i=0}^{k}t_{i}\cdot v_{i},\qquad\text{where }s+r+\sum t_{i}=1,\;s,r,t_{i}\geq 0.$
The quotient map $q$ identifies points in $\bar{\sigma}$ via the relation
$s\cdot a+\sum_{i=0}^{k}t_{i}\cdot v_{i}\approx(s+t_{0})\cdot a+\sum_{i=1}^{k}t_{i}\cdot v_{i},$
which encodes the strong collapse $K\smallsetminus w{\searrow\searrow\hskip 3.0pt}(K\smallsetminus w)\smallsetminus v$. that replaces the vertex $v_{0}=v$ by its dominator $a$ on the subcomplex The resulting quotient space $q(\bar{\sigma})$ corresponds to the internal collapse $w\ast{\mathrm{lk}}^{\downarrow}(v,K){\hskip 2.0pt\searrow\hskip 3.0pt}{\mathrm{lk}}^{\downarrow}(v,K)\subseteq\partial\bar{\sigma}$. Since we are identifying a contractible subcomplex with its image under a strong deformation retract, and this collapse occurs entirely within the boundary of $\bar{\sigma}$, the quotient $q(\bar{\sigma})$ remains homeomorphic to a closed $(k+2)$-ball $D^{k+2}$. Furthermore, the interior of $\bar{\sigma}$ is disjoint from the identified region and is mapped injectively to its image in $Z$. See [13, Prop. 3.4], [30, Ch. 4].
We define the corresponding cell $e_{\sigma}\subseteq Z$ to be the image of the interior of $\bar{\sigma}$, with closure $q(\bar{\sigma})\cong D^{k+2}$. The attaching map of $e_{\sigma}$ is then induced from the original attaching map $\varphi_{\sigma}\colon\partial\Delta^{k+2}\to K$ by composition with the quotient map. This yields a continuous map
$\bar{\varphi}_{\sigma}\colon\Delta^{k+2}/_{\approx}\;\to Z,$
where the quotient space $\Delta^{k+2}/_{\approx}$ is homeomorphic to $S^{k+1}$ and $\bar{\varphi}_{\sigma}$ is a homeomorphism onto its image. This defines a regular attaching map for the cell $e_{\sigma}$.
[…]
@@
$K\smallsetminus w=K_{0}{\searrow\searrow^{e}\hskip 5.0pt}K_{1}\searrow\dots\searrow\searrow^{e}\hskip 5.0ptK_{n}=L.$
Inductively, we have a corresponding sequence of pushouts:
[…]
Here, $\bar{i}\colon K_{n-1}\to Z_{n-1}$ is an inclusion, and the same reasoning applied to $K$ extends to $Z_{n-1}$. By induction, the space $Z_{n-1}$ carries a regular CW-structure and is homotopy equivalent to $K$. $Z_{n}$ is obtained after performing an elementary strong collapse to $K_{n-1}$. To complete the inductive step, we must ensure that the new attaching maps are still a homeomorphism onto its image. The only scenario that could in principle compromise regularity would be if, for some simplex $\sigma\in{\mathrm{st}}(w,K)$, the entire boundary $\partial\sigma$ were identified to a single point in $Z_{n}$ due to prior identifications. However, this situation cannot arise: since $w$ is always a vertex of $\sigma$, and since strong collapses remove only dominated vertices, the vertex $w$ is never removed. Moreover, even if all other vertices of $\sigma$ are eventually collapsed, the image of the boundary of $\sigma$ must contain at least two distinct points — namely, $w$ and one vertex from $L$ not dominated during the collapse sequence. Hence, the attaching map remains a homeomorphism onto its image.
Therefore, the final space $Z$ is a regular CW-complex homotopy equivalent to $K$. The fact that this homotopy equivalence is an $(N+1)$-deformation follows from the observation that this process was obtained in particular by internal collapses, and by [15, Prop. 1.3], such collapses induce $(N+1)$-deformations. ∎
Proposition 3.4 (Strong internal collapse).
Let $K$ be a simplicial complex of dimension $N$. If a full subcomplex $K_{0}\subseteq K$ strong collapses to a subcomplex $L_{0}$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0ptL$, where $L$ is a regular CW-complex obtained from $L_{0}$ by attaching one $k$-cell for each $k$-simplex in $K\smallsetminus K_{0}$.
@@
Follows by induction on the vertices in $V(K)\smallsetminus V(K_{0})$, using Lemma 3.3. ∎
Under the assumptions of Proposition 3.4, we say that there is a strong internal collapse from $K$ to $L$.
@@
    where the cells $e_{i}$ are in one-to-one correspondence with the simplices in ${\mathrm{st}}^{\downarrow}(w,K)$. This correspondence preserves dimensions, and the attaching maps of the cells are determined as in Lemma 3.3, resulting in a regular CW-complex structure.
@@
Part B. follows from part A. and Lemma 3.3. ∎
Building on Proposition 3.4 and Lemma 3.6, we establish the following result, which formalize the relationship between internal strong collapses, regular CW-complexes, and discrete Morse theory.
@@
Let $K$ be a finite simplicial complex and let $g\colon V(K)\to{\mathbb{R}}$ be a real-valued map. Then, there exists a discrete Morse function $f\colon K\to{\mathbb{R}}$ on $K$ such that its critical simplices are the simplices in {st(w,K):w is a strong critical vertex
 of
g
[…]
 of
[…]
. Moreover, ${\mathrm{core}}_{g}(K)\cong{\mathrm{core}}_{f}(K)$, that is the strong internal core associated to $g$ is precisely the internal core associated to $f$.
@@
$v_{\sigma}=\arg\max\{f(v)\colon v\in\sigma\text{ is a descending dominated vertex}\},$
@@
and let $a_{\sigma}$ denote the vertex dominating $v_{\sigma}$. The pairing $M$ is defined by
$(\sigma\smallsetminus\{a_{\sigma}\},\sigma)\in M\text{ if }a_{\sigma}\in\sigma,\quad\text{or}\quad(\sigma,\{a_{\sigma}\}\cup\sigma)\in M\text{ if }a_{\sigma}\notin\sigma.$
@@
To verify that $M$ is a matching, assume for contradiction that a simplex in $K$ is matched to more than one simplex. Without loss of generality, suppose $(\sigma,\{a\}\cup\sigma)\in M$ with $v_{\sigma}$ the associated descending dominated vertex. If there exists another pair ($(\{a\}\cup\sigma,\{a^{\prime},a\}\cup\sigma)\in M$, then $a^{\prime}$ dominates a vertex $v^{\prime}\in\sigma$. Since $v^{\prime}$ must also satisfy $v^{\prime}=\arg\max\{f(v)\colon v\in\sigma$ is descending dominated}}\}} this contradicts the uniqueness of $v_{\sigma}$. Alternatively, if there exists a pair $(\sigma^{\prime},\{a\}\cup\sigma)\in M$, then $a\sigma=\{a^{\prime}\}\cup\sigma^{\prime}$ implies $\sigma=\{a^{\prime}\}\cup\tau$ and $\sigma^{\prime}=\{a\}\cup\tau$ for some simplex $\tau$. This forces $\tau$ to contain two distinct dominated vertices, $v_{\sigma}$ and $v^{\prime}$, each maximizing $f(v)$, which is again a contradiction. Therefore, $M$ is a valid matching.

### v2

@@ 1. Introduction
Figure 1. Collapse (left) vs internal collapse (right).
@@ 1.1. Motivation and related work
Our work builds on the main principles of Bestvina–Brady Morse theory [8, 9]. In that framework, Morse functions are piecewise-linear (PL) functions defined on affine complexes, the critical points are the vertices, and changes in the topology of sublevel sets are studied through the descending links of vertices. Restricting to simplicial complexes and functions defined on their vertices, our theory refines this perspective by classifying vertices according to the structure of their descending links: either as descending dominated vertices (their descending links are simplicial cones, and induce internal strong collapses) or as strong critical vertices. While Bestvina–Brady Morse theory studies local changes in the homotopy type of sublevel sets, our approach also aims at the combinatorial and global reconstruction of a reduced complex induced by these local changes.
A complementary perspective is presented in [27, 28], where the authors encode the homotopy type of a simplicial complex endowed with a discrete Morse function into a combinatorially defined flow category. The classifying space of the flow category recovers the homotopy type of the original complex. However, these categories are often large and computationally challenging, limiting their application in algorithmic homotopy reconstruction.
@@ 1.2. Main results and outline
The combinatorial structure of the strong internal core admits a computational interpretation of internal strong collapses, which in turn leads to an efficient algorithm for their computation (Appendix A, Algorithms 1 and 2). An implementation is publicly available at https://github.com/ximenafernandez/Strong-Morse-Theory. To illustrate the effectiveness of this method, we apply it to examples from the Library of Triangulations [6], a repository of challenging simplicial complexes frequently used in the study of homotopy types. The results of the reductions are presented in Tables 1 and 2.
@@
The remainder of this manuscript is organized as follows. Section 2 presents discrete Morse theory as a generalization of Whitehead’s collapses. In Section 3, we introduce a strong version of discrete Morse theory, based on strong collapses, which enables the recovery of a regular structure in the reduced CW-complex (the strong internal core). Section 4 provides a combinatorial construction of the strong internal core in terms of posets and acyclic matchings. Finally, Section 5 discusses algorithms for the simplification of simplicial complexes and presents experiments on the Library of Triangulations [6]. The Appendix details algorithms for constructing the critical poset and a random strong internal core.
@@ 2. Discrete Morse theory as generalized collapses
We denote by ${\mathrm{core}}_{f}(K)$ the CW-complex obtained from the critical simplices of $f$ in Theorem 2.3, and refer to it as the internal core of $K$ associated to $f$. It is important to note that Theorem 2.3 does not provide an explicit description of how the critical cells are attached along their boundaries in ${\mathrm{core}}_{f}(K)$. In [17, Prop. 1.3, Thm. 1.6], the author offers a recursive construction of the attaching maps for the critical cells, based on the finiteness of $K$; however, the construction is only made explicit for 2-dimensional cells.
@@
Proposition 2.5 also admits a formulation in terms of CW-complexes; see [17, Proposition 1.3]. We refer to the deformation from $K$ to $L$ as an internal collapse. The collapse takes place within a subcomplex $K_{0}\subseteq K$, while the resulting deformation affects the entire complex through the reattachment of cells corresponding to the simplices in $K\smallsetminus K_{0}$.
@@ 3. Strong discrete Morse theory
Given a simplicial complex $K$ and a vertex $v\in V(K)$, the open star of $v$, denoted ${\mathrm{st}}(v,K)$, is the collection of simplices in $K$ that contain $v$. The subcomplex of simplices disjoint from $v$ is denoted $K\smallsetminus v$. The link of $v$ in $K$, denoted ${\mathrm{lk}}(v,K)$, is the subcomplex of $K\smallsetminus v$ consisting of simplices $\sigma$ such that $\sigma\cup\{v\}$ is a simplex of $K$. A simplicial cone over $K$ with apex $a$ (a vertex not in $K$) is the complex $a*K$ consisting of all simplices of $K$, the vertex $\{a\}$, and all simplices of the form $\sigma\cup\{a\}$ with $\sigma\in K$. The closed star of $v$, denoted $\overline{{\mathrm{st}}}(v,K)$, is the subcomplex $v\ast{\mathrm{lk}}(v,K)$.
@@
Let $K$ be a simplicial complex. A vertex $v$ of $K$ is dominated by a vertex $a$ if ${\mathrm{lk}}(v,K)$ is a simplicial cone $a*K_{0}$, where $K_{0}$ is a subcomplex of $K$. In that case, there is an elementary strong collapse from $K$ to $K\smallsetminus v$, denoted $K\searrow K\smallsetminus v$. A strong collapse $K{\searrow\searrow\hskip 3.0pt}L$ from a simplicial complex $K$ to a subcomplex $L$ is a sequence of elementary strong collapses starting at $K$ and ending at $L$. In that case, we say that $L$ is a strong core of $K$. If $L$ has no dominated vertices, we say that $L$ is a minimal strong core of $K$ [^2] This is simply called a core in [3]. We will use the term minimal strong core in this manuscript to avoid confusion with other core constructions.
If $v\in K$ is dominated by $a$, then there is an explicit strong deformation retraction $|r_{a}|\colon|K|\to|K\smallsetminus v|$ induced by the simplicial map
(1) $r_{a}(\sigma)=\begin{cases}\sigma&\text{if }v\notin\sigma,\\ \{a\}\cup\sigma\smallsetminus\{v\}&\text{if }v\in\sigma.\end{cases}$
@@
with $K_{i+1}=K_{i}\smallsetminus v_{i}$ and $v_{i}$ dominated by $a_{i}\in K_{i}$, then there is a strong deformation retraction $|r|\colon|K|\to|L|$ induced by the composition
(2) $r:=r_{a_{n}}\circ r_{a_{n-1}}\circ\dots\circ r_{a_{1}}.$
@@
Recall that a CW-complex $K$ is regular if the attaching map $\varphi\colon S^{n-1}\to K^{(n-1)}$ of each open $n$-cell $e^{n}$ is a homeomorphism onto its image $\partial e^{n}$ in the $(n-1)$-skeleton. A regular CW-complex is said to be combinatorial if, for each cell $e^{n}$, the attaching map — regarded as a cellular map from $S^{n-1}$ endowed with a CW-structure — is a combinatorial map, meaning that it maps each open $k$-cell of $S^{n-1}$ homeomorphically onto an open $k$-cell of $K$ (see [21, Ch. II]). The geometric realization $|K|$ of any simplicial complex $K$ is canonically a regular combinatorial CW-complex. Given a vertex $v$ in a regular CW-complex $K$, we define the open star ${\mathrm{st}}(v,K)$ as the union of all open cells in $K$ that contain $v$ as a face. We denote by $\overline{e}$ the closure of a cell $e\in K$, i.e., the smallest subcomplex of $K$ containing $e$, which consists of $e$ together with all its faces. Similarly, we denote by $K\smallsetminus v$ the subcomplex of $K$ consisting of all cells $e$ such that $v\notin\overline{e}$.
The following result formalizes the notion of an internal strong collapse, illustrated in Figure 2.
Theorem 3.3.
Let $K$ be a regular combinatorial CW-complex, and let $w$ be a vertex of $K$ such that the subcomplex $K\smallsetminus w$ is a simplicial complex. Suppose there exists a subcomplex $L\subseteq K\smallsetminus w$ such that $K\smallsetminus w{\searrow\searrow\hskip 3.0pt}L$ via a sequence of strong collapses.
Then there exists a regular combinatorial CW-complex
$Z=L\cup\bigcup_{\ell}\tilde{e}_{\ell}$
that is homotopy equivalent to $K$, where:
  • $L$ is embedded as a subcomplex of $Z$,
  • for each cell $e_{\ell}\in K$ such that $w\in\overline{e_{\ell}}$, there is a corresponding $\ell$-cell $\tilde{e}_{\ell}$ in $Z$ not contained in $L$, and
  • the attaching map of each $\tilde{e}_{\ell}$ is induced from that of $e_{\ell}$ and the strong deformation retraction $|r|\colon|K\smallsetminus w|\to|L|$ determined by the strong collapse.
Moreover, if $\dim(K)=N$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0ptZ$.
@@
where $i$ is the inclusion map and $|r|\colon|K\smallsetminus w|\to|L|$ is the strong deformation retraction induced by the strong collapse. Since $|r|$ is a homotopy equivalence and $i$ is a closed cofibration, the Gluing Theorem (see, e.g., [13, 7.5.7, Corollary 2]) implies that $\bar{r}$ is also a homotopy equivalence.
We now define a regular combinatorial CW-structure on $Z$. Assume first that the collapse $K\smallsetminus w{\searrow\searrow\hskip 3.0pt}L$ is elementary, i.e., $L=(K\smallsetminus w)\smallsetminus v$ for some vertex $v$ dominated by $a\in K\smallsetminus w$.
The 00-skeleton of $Z$ is defined as $Z^{(0)}=L^{(0)}\cup\{w\}$. We define a map $q\colon K^{(0)}\to Z^{(0)}$ by
$q(x)=\begin{cases}x&\text{if }x\in L^{(0)}\cup\{w\},\\ a&\text{if }x=v,\end{cases}$
which agrees with the retraction $|r|$ on $K^{(0)}\smallsetminus\{w\}$.
Inductively, suppose that a regular combinatorial CW-structure has been defined on $Z^{(\ell-1)}$, and that the map $q\colon K^{(\ell-1)}\to Z^{(\ell-1)}$ is combinatorial and coincides with the simplicial retraction $|r|$ on $(K\smallsetminus w)^{(\ell-1)}$. Let $e$ be an $\ell$-cell in $K$ with attaching map $\varphi\colon S^{\ell-1}\to K^{(\ell-1)}$.
If $w\notin\overline{e}$, then $e$ belongs entirely to the subcomplex $K\smallsetminus w$, and its image in $Z$ is defined via $q=|r|$. Since $r\colon K\smallsetminus w\to L$ is simplicial, and the attaching maps in $K\smallsetminus w$ are simplicial, it follows that $q(e)$ inherits a regular combinatorial CW-structure. In particular, if $e\in L$, then $q$ acts as the identity on $e$.
If $w\in\overline{e}$, we distinguish three cases:
  1. (1)
    If $v\notin\overline{e}$, then $q$ acts as the identity on $e$, and the attaching map of $e$ remains unchanged in $Z$, inheriting the regular combinatorial CW-structure from $K$.
  2. (2)
    If $v\in\overline{e}$ but $a\notin\overline{e}$, then $r$ sends $v$ to $a$ in $\partial e\cap(K\smallsetminus w)$, resulting in a relabelling, inducing a bijective identification on the boundary. The attaching map $\tilde{\varphi}\colon S^{\ell-1}\to Z^{(\ell-1)}$ of $\tilde{e}\in Z$ is given by $\tilde{\varphi}:=r\circ\varphi$, and the image cell $q(e)$ is obtained from $e$ via this relabelling.
  3. (3)
    If both $v$ and $a$ belong to $\overline{e}$, then the retraction $r$ modifies the attaching map non-trivially. Since $K$ is regular and combinatorial, the boundary $\partial e$ is a PL $(\ell-1)$-sphere. Let
    $A:=|\overline{{\mathrm{st}}}(v,K\smallsetminus w)|\cap\partial e,\quad B:=|{\mathrm{lk}}(v,K\smallsetminus w)|\cap\partial e.$
    Then $A$ is a PL $(\ell-1)$-ball and $B$ a PL $(\ell-2)$-ball, and the strong collapse induces a retraction $r|_{A}\colon A\to B$. By [32, Ch. 3], there exist regular neighbourhoods $U\supseteq A$ and $V\supseteq B$ in $\partial e$ and a map $f\colon\partial e\to\partial e$ such that $f|_{A}=r|_{A}$, and $f|_{\partial e\setminus A}\colon\partial e\setminus A\to\partial e\setminus B$ is a homeomorphism. This induces a homeomorphism
    $\bar{f}\colon\partial e/\!\sim\ \longrightarrow\partial e,\quad x\sim r(x)\ \text{for }x\in A.$
    If $\varphi\colon S^{\ell-1}\to\partial e$ is the original attaching map of $e$, then the modified attaching map is given by:
    ${S^{\ell-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${\partial e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${S^{\ell-1}/\!\approx\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varphi}}$${q(\partial e)\subseteq Z^{(\ell-1)}}$
    where $x\approx y$ if $\varphi(x)\sim\varphi(y)$ under the identification induced by $r$. Since $\bar{f}$ is a homeomorphism, the quotient space $S^{\ell-1}/\!\approx$ is homeomorphic to $S^{\ell-1}$, and the new attaching map $\tilde{\varphi}$ is a homeomorphism onto its image. Therefore, the cell $\tilde{e}$ is attached regularly, and inherits a combinatorial CW-structure. Define $q(e)=\tilde{e}$.
@@
$K\smallsetminus w=K_{0}{\searrow\searrow\hskip 3.0pt}K_{1}{\searrow\searrow\hskip 3.0pt}\cdots{\searrow\searrow\hskip 3.0pt}K_{n}=L.$
At each step, the same argument applies to the retraction $K_{i}\to K_{i+1}$, ensuring that the regularity and combinatorial structure are preserved. Since $w$ is not removed in the process, the identification maps never collapse an entire boundary sphere to a point.
Thus, $Z$ is a regular combinatorial CW-complex homotopy equivalent to $K$. The fact that this is an $(N+1)$-deformation follows from [17, Prop. 1.3]. ∎
The previous result generalizes to the case where $K_{0}{\searrow\searrow\hskip 3.0pt}L_{0}$ and $K_{0}$ is obtained from a simplicial complex $K$ by successively removing vertices (together with their open stars). In this case, $K_{0}$ is a full subcomplex of $K$, meaning that every simplex of $K$ whose vertices all belong to $V(K_{0})$ is itself a simplex in $K_{0}$.
Theorem 3.4 (Strong internal collapse).
Let $K$ be a simplicial complex of dimension $N$, and let $K_{0}\subseteq K$ be a full subcomplex. Suppose $K_{0}$ strong collapses to a subcomplex $L_{0}$. Then $K$ is $(N+1)$-homotopy equivalent to a combinatorial regular CW-complex $L$ obtained from $L_{0}$ by attaching one $k$-cell for each $k$-simplex in $K\smallsetminus K_{0}$.
@@
Follows by induction on the vertices in $V(K)\smallsetminus V(K_{0})$, applying Theorem 3.3 at each step. ∎
Under the assumptions of Theorem 3.4, we say that there is a strong internal collapse from $K$ to $L$.
@@
    where the cells $e_{i}$ are in one-to-one correspondence with the simplices in ${\mathrm{st}}^{\downarrow}(w,K)$. This correspondence preserves dimensions, and the attaching maps of the cells are determined as in Theorem 3.3, resulting in a regular CW-complex structure.
@@
Part B. follows from part A. and Theorem 3.3. ∎
Building on Theorem 3.4 and Lemma 3.6, we establish the following result, which formalize the relationship between internal strong collapses, regular CW-complexes, and discrete Morse theory.
@@
Let $K$ be a finite simplicial complex and let $g\colon V(K)\to{\mathbb{R}}$ be a real-valued map. Then, there exists a discrete Morse function $f\colon K\to{\mathbb{R}}$ on $K$ such that its critical simplices are the simplices in $\{{\mathrm{st}}^{\downarrow}(w,K):w\text{ is a strong critical vertex of }g\}$. Moreover, ${\mathrm{core}}_{g}(K)\cong{\mathrm{core}}_{f}(K)$, that is the strong internal core associated to $g$ is precisely the internal core associated to $f$.
@@
$v_{\sigma}=\arg\max\{g(v)\colon v\in\sigma\text{ is a descending dominated vertex}\},$
@@
and let $a_{\sigma}$ denote the vertex dominating $v_{\sigma}$. The pairing $M$ is then defined as follows: if $a_{\sigma}\in\sigma$, we include the pair $(\sigma\setminus\{a_{\sigma}\},\sigma)$ in $M$; otherwise, we include the pair $(\sigma,\sigma\cup\{a_{\sigma}\})$ in $M$.
@@
To verify that $M$ is a matching, assume for contradiction that a simplex in $K$ is matched to more than one simplex. Without loss of generality, suppose $(\sigma,\{a\}\cup\sigma)\in M$ with $v_{\sigma}$ the associated descending dominated vertex. If there exists another pair ($(\{a\}\cup\sigma,\{a^{\prime},a\}\cup\sigma)\in M$, then $a^{\prime}$ dominates a vertex $v^{\prime}\in\sigma$. Since $v^{\prime}$ must also satisfy $v^{\prime}=\arg\max\{g(v)\colon v\in\sigma$ is descending dominated}}\}} this contradicts the uniqueness of $v_{\sigma}$. Alternatively, if there exists a pair $(\sigma^{\prime},\{a\}\cup\sigma)\in M$, then $a\sigma=\{a^{\prime}\}\cup\sigma^{\prime}$ implies $\sigma=\{a^{\prime}\}\cup\tau$ and $\sigma^{\prime}=\{a\}\cup\tau$ for some simplex $\tau$. This forces $\tau$ to contain two distinct dominated vertices, $v_{\sigma}$ and $v^{\prime}$, each maximizing $g(v)$, which is again a contradiction. Therefore, $M$ is a valid matching.